### v1

# Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach

Abstract.
Decreasing costs and new technologies have led to an increase in the amount of data available to plant breeding programs. High-throughput phenotyping (HTP) platforms routinely generate high-dimensional datasets of secondary features that may be used to improve genomic prediction accuracy. However, integration of this data comes with challenges such as multicollinearity, parameter estimation in $p>n$ settings, and the computational complexity of many standard approaches. Several methods have emerged to analyze such data, but interpretation of model parameters often remains challenging.
We propose genetic factor best linear unbiased prediction (gfBLUP), a seven-step prediction pipeline that reduces the dimensionality of the original secondary HTP data using generative factor analysis. In short, gfBLUP uses redundancy filtered and regularized genetic and residual correlation matrices to fit a maximum likelihood factor model and estimate genetic latent factor scores. These latent factors are subsequently used in multi-trait genomic prediction. Our approach performs on par or better than alternatives in extensive simulations and a real-world application, while producing easily interpretable and biologically relevant parameters. We discuss several possible extensions and highlight gfBLUP as the basis for a flexible and modular multi-trait genomic prediction framework.
Key words: Empirical Bayes; Factor analysis; Genomic prediction; High-dimensional data
1. Introduction
Genomic prediction has revolutionized plant breeding programs over the past two decades. Its use allows maintaining and evaluating breeding populations larger than ever before. As a result, breeding programs have seen increases in efficiency leading to more rapid development of varieties that meet todays challenges [1]. Genomic prediction is a powerful tool in ensuring food security in the face of a growing population and more extreme climate conditions [2].
The practice of genomic prediction involves training a statistical model using evaluated genotypes, followed by making trait predictions for untrialed test genotypes [3]. The model is mainly used in a univariate form (gBLUP, genomic best linear unbiased prediction) where the predictions are sums of genomic marker effects [4]. These predictions allow breeders to make selections without the need for extensive field trials.
Recent years have seen an increase in the use of high-throughput phenotyping (HTP) technologies resulting in the characterization of individuals in more aspects than just those of the genomic and trait phenotypic data. Examples include the collection of hyperspectral reflectivity measurements on the crop canopy in large field trials, metabolic profiling of plant leaves, and gene expression profiling using DNA microarrays [e.g., 5]. These secondary phenotypes can be collected for either the training set alone (termed CV1, cross-validation scenario 1), or both the training and test individuals (CV2, cross-validation scenario 2) [6]. If of sufficient quality, the secondary data can contain valuable information that readily improves genomic predictions for the trait of interest (focal trait). There are several such examples where integration of genomic and phenomic data leads to higher accuracies [7]
Univariate genomic prediction uses only genomic information and focal trait information to obtain predictions. To integrate phenomics and genomics, the univariate model can be extended to a multivariate form relatively easily by modelling the focal trait as well as the secondary phenotypes as functions of the SNP markers [8]. Direct integration of HTP data to improve predictions is challenging for several reasons, however. First of all, secondary HTP data is often high-dimensional complicating parameter estimation. Second, subsets of secondary features are often highly correlated. This multicollinearity can greatly complicate the interpretation of model parameters [9]. Finally, even if the number of observations were greater than the number of features, and features were not highly correlated, direct integration would still be computationally demanding as the calculation of the BLUPs requires inverting a symmetric matrix whose dimension equals the number of genotypes times the number of features.
1.1. Current methodology
A number of methods have been developed to deal with the high dimensionality, multicollinearity, and computational complexity. Lopez-Cruz et al. [11] proposed linearly combining the secondary features into a single regularized selection index (siBLUP), thereby reducing the dimensionality to one. The lsBLUP method introduced by Arouisse et al. [8] is similar in the sense that it reduces the dimensionality to one, but uses a LASSO prediction of the focal trait using the HTP features instead of a regularized selection index. Another alternative is MegaLMM as proposed by Runcie et al. [10]. MegaLMM constructs an arbitrary number of independent factors and errors that are linear combinations of the original features. Finally, neural networks and deep learning have also been investigated as options for the integration of secondary HTP and genomic data [12]. Due to their flexible nature they present an attractive option to model complex interactions between many features.
1.2. The proposed model & overview
We introduce gfBLUP (genetic factor BLUP) to deal with the challenges of secondary HTP data. It is based on the idea that HTP features typically form a set of many noisy measurements of a much lower dimensional set of latent features [13]. In short, we use factor analysis to reduce the dimensionality of the HTP data to a data-driven number of uncorrelated latent factors which are subsequently used in multivariate genomic prediction. We avoid any manual hyperparameter tuning by using a data-driven approach to determine the dimensionality of the factor model. Our approach is also flexible in the sense that it can accommodate latent structures of any dimension lower than the well-known Ledermann bound [14]. This flexibility means that complex structures underlying high-dimensional data can be accurately modeled. Furthermore, the factor loadings and genetic correlations between factors and the focal trait are interpretable and reveal the relevant biological patterns in the otherwise noisy HTP data.
The remainder of this paper is organized as follows. We elaborate the proposed factor analytic and multivariate genomic prediction approach in Section 2, showing how a seven-step prediction pipeline produces predicted breeding values as well as interpretable parameters. Section 3 provides details and results on a comprehensive simulation study comparing the performances of different methods as well as the importance of various genetic parameters. In Section 4 we apply gfBLUP and alternatives to a wheat yield and hyperspectral reflectivities dataset to evaluate their real-world performance and show how gfBLUP produces biologically relevant and interpretable parameters. Finally, we discuss several possible extensions to the proposed approach in Section 5. Information on notation can be found in Section S.1 of the Supplementary Material (SM).
2. Model
2.1. Model and assumptions
Consider the full phenotypic data matrix $\mathbf{Y}_{sf}\in\mathbb{R}^{n\times(p+1)}$ of the secondary features and focal trait. We assume a balanced design for simplicity, but note that this is not required. In a balanced design $n=n_{g}r$, i.e., the number of observations is equal to the number of genotypes multiplied by the number of replicates per genotype. The matrix $\mathbf{Y}_{sf}$ can be partitioned:
(1) $\mathbf{Y}_{sf}=\begin{bmatrix}\mathbf{Y}_{s}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
where the matrix $\mathbf{Y}_{s}$ is the $n\times p$ matrix containing the mean-centered secondary features and $\mathbf{y}_{f}$ is the $n\times 1$ column vector containing the mean-centered focal trait. Assuming no covariates, all phenotypes are composed of genetic and residual components:
$\mathbf{Y}_{sf}=\mathbf{G}_{sf}+\mathbf{E}_{sf}\text{.}$
We assume that the genetic components are independent of the residual components so that
$\operatorname{vec}({\mathbf{Y}_{sf}})=\operatorname{vec}({\mathbf{G}_{sf}})+\operatorname{vec}({\mathbf{E}_{sf}})\sim\mathcal{N}_{n(p+1)}(\mathbf{0},\bm{\Sigma}^{g}\otimes\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}+\bm{\Sigma}^{\epsilon}\otimes\mathbf{I}_{n})\text{,}$
where $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}$ and $\mathbf{I}_{n}$ are the row covariances of $\mathbf{G}_{sf}$ and $\mathbf{E}_{sf}$, respectively. The $n\times n_{g}$ incidence matrix $\mathbf{Z}$ links individuals to genotypes. Similarly, $\bm{\Sigma}^{g}$ and $\bm{\Sigma}^{\epsilon}$ are the column covariances of $\mathbf{G}_{sf}$ and $\mathbf{E}_{sf}$. These covariance matrices can be block-partitioned:
$\begin{split}\bm{\Sigma}^{g}&=\begin{pmatrix}\bm{\Sigma}_{ss}^{g}&\bm{\sigma}_{sf}^{g}\\ \bm{\sigma}_{fs}^{g}&\sigma_{ff}^{g}\\ \end{pmatrix}=\begin{pmatrix}\bm{\Sigma}_{s.}^{g}\\ \bm{\sigma}_{f.}^{g}\end{pmatrix}\\ \bm{\Sigma}^{\epsilon}&=\begin{pmatrix}\bm{\Sigma}_{ss}^{\epsilon}&\bm{\sigma}_{sf}^{\epsilon}\\ \bm{\sigma}_{fs}^{\epsilon}&\sigma_{ff}^{\epsilon}\\ \end{pmatrix}=\begin{pmatrix}\bm{\Sigma}_{s.}^{\epsilon}\\ \bm{\sigma}_{f.}^{\epsilon}\end{pmatrix}\text{,}\end{split}$
where $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$ contain the genetic and residual covariances among secondary features, $\sigma_{ff}^{g}$ and $\sigma_{ff}^{\epsilon}$ are the scalar focal trait genetic and residual variances, and $\bm{\sigma}_{sf}^{g}=(\bm{\sigma}_{fs}^{g})^{\top}$ and $\bm{\sigma}_{sf}^{\epsilon}=(\bm{\sigma}_{fs}^{\epsilon})^{\top}$ contain the genetic and residual covariances between secondary features and the focal trait. The method we propose is initially unsupervised as we first focus on the secondary feature data matrix $\mathbf{Y}_{s}$ and the corresponding covariance matrices $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$. We then use focal trait training data to train the prediction model.
While we require plot-level data to estimate covariance matrices, genomic prediction typically uses best linear unbiased estimates (BLUEs). For a completely randomized design we can obtain these BLUEs for a single genotype $j$ by simply taking the genotypic mean of all replicates $i=1,\dots,r$ [15]:
(2) $\bar{\mathbf{y}}_{s(j)}=\mathbf{g}_{s(j)}+\bar{\bm{\epsilon}}_{s(j)}=\mathbf{g}_{s(j)}+\dfrac{1}{r}\sum_{i=1}^{r}\bm{\epsilon}_{s(i|j)}\text{.}$
Note that as the residual vectors $\bm{\epsilon}_{s(i|j)}$ of the replicates of genotype $j$ are iid random vectors, basic covariance algebra shows that the covariance of $\bar{\bm{\epsilon}}_{s(j)}$ is simply $r^{-1}\bm{\Sigma}_{ss}^{\epsilon}$ [15] (see also SM Section S.3). As a result, $\bar{\mathbf{y}}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma}_{ss}^{p})$, where $\bm{\Sigma}_{ss}^{p}=\bm{\Sigma}_{ss}^{g}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon}$, assuming for a moment that $\mathbf{K}\circ\mathbf{I}_{n_{g}}=\mathbf{I}_{n_{g}}$.
In gfBLUP, we use a factor analytic structure for $\bm{\Sigma}_{ss}^{g}$. This follows from the assumption that the genetic components $\mathbf{g}_{s(j)}$ can be written as a weighted sum of $m$ latent factors and $p$ errors:
(3) $\mathbf{g}_{s(j)}^{\top}=\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j}\text{,}$
where $\mathbf{\Lambda}$ is a $p\times m$ matrix containing factor loadings, $\bm{\xi}_{j}$ is an $m\times 1$ vector containing the latent factor scores for genotype $j$, and $\bm{\varepsilon}_{j}$ contains the factor model errors. Substituting (3) into (2) gives us the full model for the phenotypic data:
(4) $\bar{\mathbf{y}}_{s(j)}^{\top}=\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j}+\bar{\bm{\epsilon}}_{s(j)}^{\top}\text{.}$
We now make the following assumptions: (i) $\text{rank}(\mathbf{\Lambda})=m<p$, (ii) $\bm{\varepsilon}_{j}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{\Psi})$, where $\mathbf{\Psi}=\mathbb{E}(\bm{\varepsilon}_{j}\bm{\varepsilon}_{j}^{\top})$ and $\mathbf{\Psi}\equiv diag(\psi_{11},\dots,\psi_{pp})$ with $\psi_{qq}>0,\forall qq$, (iii) $\bm{\xi}_{j}\sim\mathcal{N}_{m}(\mathbf{0},\mathbf{I}_{m})$, (iv) $\bm{\xi}_{j}\ConIndepNat\bm{\varepsilon}_{j^{\prime}},\forall j,j^{\prime}$, (v) $\bm{\xi}_{j}\ConIndepNat\bar{\bm{\epsilon}}_{s(j^{\prime})},\forall j,j^{\prime}$, and (vi) $\bm{\varepsilon}_{j}\ConIndepNat\bar{\bm{\epsilon}}_{s(j^{\prime})},\forall j,j^{\prime}$.
These assumptions imply that the genetic covariance matrix $\bm{\Sigma}_{ss}^{g}$ equals the model-implied genetic covariance matrix $\bm{\Sigma}_{ss}^{g}(\bm{\Theta})=\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}=\mathbb{E}\left[(\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j})(\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j})^{\top}\right]$ where $\bm{\Theta}=\{m,\mathbf{\Lambda},\mathbf{\Psi}\}$. Therefore, $\bar{\mathbf{y}}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon})$. The covariance of the genetic components $\mathbf{g}_{s(j)}$ can thus be decomposed into common ($\mathbf{\Lambda}\mathbf{\Lambda}^{\top}$) and unique ($\mathbf{\Psi}$) elements. Using standard covariance algebra we can also obtain the joint distribution of the secondary feature BLUEs and latent factors as
(5) $\begin{bmatrix}\bar{\mathbf{y}}_{s(j)}^{\top}\\ \bm{\xi}_{j}\end{bmatrix}\sim\mathcal{N}_{p+m}\left[\mathbf{0},\bm{\Sigma}(\bm{\Theta})_{\xi}^{y}\right]\text{,}$
where
$\bm{\Sigma}(\bm{\Theta})_{\xi}^{y}=\begin{bmatrix}\bm{\Sigma}(\bm{\Theta})_{yy}&\bm{\Sigma}(\bm{\Theta})_{y\xi}\\ \bm{\Sigma}(\bm{\Theta})_{\xi y}&\bm{\Sigma}(\bm{\Theta})_{\xi\xi}\end{bmatrix}=\begin{bmatrix}\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon}&\mathbf{\Lambda}\\ \mathbf{\Lambda}^{\top}&\mathbf{I}_{m}\end{bmatrix}=\begin{bmatrix}\bm{\Sigma}_{ss}^{g}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon}&\mathbf{\Lambda}\\ \mathbf{\Lambda}^{\top}&\mathbf{I}_{m}\end{bmatrix}\text{.}$
We will use the model above to project the observed secondary features onto a lower-dimensional latent space, followed by multivariate genomic prediction in seven steps. We first use the plot-level data to obtain estimates for the genetic and residual correlation matrices of the observed features (I). We then redundancy filter the data (II) to ensure there are no pairs of features with near perfect genetic correlations, followed by regularization of the filtered genetic and residual correlation matrices (III). We then estimate the latent dimension (IV) and fit the factor model (V). Next, we obtain estimates for the latent factor scores (VI) and use those to select a subset and calculate multivariate genomic predictions using the standard BLUP equations (VII). The remainder of this section is devoted to the details of these seven steps.
2.2. Covariance matrix estimation (step I)
The factor analytic approach in gfBLUP largely revolves around the genetic and residual covariance matrices of the observed secondary data. We can obtain estimates of these matrices by using REML (restricted maximum likelihood) and the kinship matrix $\mathbf{K}$ for a small number of features. When the number of secondary features $p$ exceeds $5\text{-}10$, however, the computational complexity becomes prohibitive and REML estimates can no longer be obtained [16]. As this is usually the case for high-dimensional HTP data, we use data containing genotypic replicates and correct for any experimental design factors to obtain pseudo-CRD (pseudo completely randomized design) data as a pre-processing step. Note again that in this case, assuming balanced data, the total number of samples is simply the number of genotypes times the number of replicates, i.e., $n=n_{g}r$. We then assume that $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}=\mathbf{Z}\mathbf{Z}^{\top}$, i.e., $\mathbf{K}=\mathbf{I}_{n_{g}}$, and obtain estimates for $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$ using sums of squares as measures for between and within-genotype variation [17]:
(6) $\begin{split}\mathbf{MS}_{ss}^{g}&=(\bar{\mathbf{Y}}_{s}-\bar{\mathbf{Y}}_{s}^{g})^{\top}(\bar{\mathbf{Y}}_{s}-\bar{\mathbf{Y}}_{s}^{g})(n_{g}-1)^{-1}\\ \mathbf{MS}_{ss}^{\epsilon}&=(\bar{\mathbf{Y}}_{s}^{g}-\mathbf{Y}_{s})^{\top}(\bar{\mathbf{Y}}_{s}^{g}-\mathbf{Y}_{s})(n-n_{g})^{-1}\\ \hat{\bm{\Sigma}}_{ss}^{g}&=(\mathbf{MS}_{ss}^{g}-\mathbf{MS}_{ss}^{\epsilon})r^{-1}\\ \hat{\bm{\Sigma}}_{ss}^{\epsilon}&=\mathbf{MS}_{ss}^{\epsilon}\text{,}\end{split}$
where $\mathbf{Y}_{s}$ is the $n\times p$ data matrix containing the secondary features. Overall means and genotypic means are located in $\bar{\mathbf{Y}}_{s}$ and $\bar{\mathbf{Y}}_{s}^{g}$, respectively. A well-known issue with the estimation of such covariance matrices using multivariate sums of squares is the fact that sufficient within-genotype variation can result in $\hat{\bm{\Sigma}}_{ss}^{g}$ having negative eigenvalues [18] as it is not a moment matrix but rather the difference of two moment matrices (as opposed to $\hat{\bm{\Sigma}}_{ss}^{\epsilon}$ which is a moment matrix and therefore at worst positive-semidefinite if $p>n$). We deal with this issue by simply finding the nearest positive-definite matrix using Higham’s algorithm [19]. From here on, $\hat{\bm{\Sigma}}_{ss}^{g}$ refers to this nearest positive definite matrix.
We scale the covariance matrices as our approach requires them to be on the correlation scale for regularization:
$\begin{split}\hat{\mathbf{R}}_{ss}^{g}&=(\hat{\bm{\Sigma}}_{ss}^{g}\circ\mathbf{I}_{p})^{-1/2}\hat{\bm{\Sigma}}_{ss}^{g}(\hat{\bm{\Sigma}}_{ss}^{g}\circ\mathbf{I}_{p})^{-1/2}\\ \hat{\mathbf{R}}_{ss}^{\epsilon}&=(\hat{\bm{\Sigma}}_{ss}^{\epsilon}\circ\mathbf{I}_{p})^{-1/2}\hat{\bm{\Sigma}}_{ss}^{\epsilon}(\hat{\bm{\Sigma}}_{ss}^{\epsilon}\circ\mathbf{I}_{p})^{-1/2}\text{.}\end{split}$
We emphasize that the assumption of independent genetic effects is only made for the genetic and residual covariance estimators in equation (6). We use a marker-based kinship matrix in the final genomic prediction step. The assumption of diagonal kinship ($\mathbf{K}=\mathbf{I}_{n_{g}}$) in (6) only means that the subsequent factor analysis is based on the covariance between the total (rather than only the additive) genetic effects. This seems a reasonable approximation since for most quantitative traits the genetic variance is predominantly additive [20].
2.3. Redundancy filtering (step II)
Given the fact that multicollinearity is common in HTP data, some of the $p$ secondary features will likely show near perfect genetic correlation and are therefore redundant. We filter the genetic correlation matrix using the algorithm described by Peeters et al. [21] with an absolute genetic correlation threshold $\tau$ (i.e., after filtering, no two features will have an absolute genetic correlation $|\rho^{g}|$ equal to, or exceeding $\tau$). This filtering step produces $\hat{\mathbf{R}}_{ss}^{g*}\in\mathbb{R}^{p^{*}\times p^{*}}$ containing only the $p^{*}$ secondary features that pass filtering ($p^{*}\leq p$). We subset the residual correlation matrix and data matrix so they too only contain the remaining $p^{*}$ features.
2.4. Regularization (step III)
The filtered estimates $\hat{\mathbf{R}}_{ss}^{g*}$ and $\hat{\mathbf{R}}_{ss}^{\epsilon*}$ may still be ill-conditioned or singular which is why regularization is required. To achieve this, we use penalized maximum likelihood forms in the remainder of the gfBLUP method [21, 22]:
(7) $\begin{split}\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g})&=(1-\vartheta_{g})\hat{\mathbf{R}}_{ss}^{g*}+\vartheta_{g}\mathbf{I}_{p^{*}}\\ \hat{\mathbf{R}}_{ss}^{\epsilon*}(\vartheta_{\epsilon})&=(1-\vartheta_{\epsilon})\hat{\mathbf{R}}_{ss}^{\epsilon*}+\vartheta_{\epsilon}\mathbf{I}_{p^{*}}\text{,}\end{split}$
where $\vartheta_{g},\vartheta_{\epsilon}\in(0,1]$. We find the optimal values for the penalty parameters $\vartheta_{g}^{\ddagger}$ and $\vartheta_{\epsilon}^{\ddagger}$ through $K$-fold cross-validation where we split the data by dividing the $n_{g}$ genotypes over $K$ disjoint folds and minimize the following expression using Brent’s algorithm [23]:
(8) $\varphi^{K}(\vartheta):=\dfrac{1}{K}\sum_{k=1}^{K}w_{k}\left\{\text{ln}\Big{|}\Bigl{(}\mathbf{R}_{ss}^{*}(\vartheta)\Bigr{)}_{\neg k}\Big{|}+\operatorname{tr}\left[\Bigl{(}\mathbf{R}_{ss}^{*}\Bigr{)}_{k}\Bigl{(}\mathbf{R}_{ss}^{*}(\vartheta)\Bigr{)}_{\neg k}^{-1}\right]\right\}\text{,}$
where $\left(\mathbf{R}_{ss}^{*}(\vartheta)\right)_{\neg k}$ is the penalized genetic or residual correlation matrix given the corresponding penalty $\vartheta$ estimated using the $K-1$ folds. The non-penalized correlation matrix $(\mathbf{R}_{ss}^{*})_{k}$ is estimated using the left-out fold. Each left-out fold is weighed by $w_{k}$, which is equal to the number of genotypes or individuals in that fold for the genetic and residual correlation matrix, respectively. See Peeters et al. [21] for further details. We use an efficient implementation to minimize (8) which results in iterations of complexity $O(p^{*})$ as opposed to $O(p^{*3})$ due to the repeated inversion of $\mathbf{R}_{ss}^{*}(\vartheta)_{\neg k}$ required in the naive minimization. This leads to faster optimization, especially if optimal penalty values are close to the bounds. Details regarding our implementation can be found in SM Section S.4. Note that both $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$ and $\hat{\mathbf{R}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})$ balance the unbiased but high-variance estimated correlation matrices with the biased but stable identity matrix, optimizing the bias-variance trade-off. We may use another target matrix $\mathbf{T}$ instead of the identity if prior information is available in the form of, for example, covariance estimates from earlier experiments. The only requirement is that $\mathbf{T}$ is a positive-definite correlation matrix.
2.5. Latent dimension (step IV)
In principle, the model specified in (4) is valid for any latent dimension $m\leq\frac{1}{2}\left[2p+1-(8p+1)^{1/2}\right]$ [14]. We must specify an estimate for $m$ prior to fitting the factor model and obtaining an estimate for the other parameters in $\bm{\Theta}$, however. To do so we use a lower eigenvalue bound based on the Marchenko-Pastur distribution [24]:
$\omega_{+}=\left(1+\sqrt{\dfrac{p^{*}}{n_{g}}}\right)^{2}\text{,}$
where $n_{g}$ is the number of genotypes used to estimate the genetic covariance matrix as described in Section 2.2. We then determine $\tilde{m}$:
$\tilde{m}:=\text{card}\left\{d\left[\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})\right]>\omega_{+}\right\}\text{,}$
where $d[\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})]$ are the eigenvalues of $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$. Consequently, the latent dimension we use is simply the number of eigenvalues of the regularized correlation-scale estimate for $\bm{\Sigma}_{ss}^{g*}$, $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$, that exceed $\omega_{+}$.
2.6. Factor analytic modeling (step V)
With regularized estimates for the correlation matrices and latent dimension available, we now proceed with the factor analytic dimension reduction of the redundancy filtered secondary data. In the context of plant breeding, we are mainly interested in genetic correlations and will therefore fit the factor model using a reduced form of the secondary data covariance matrix. Recall that the full covariance of the BLUEs given in Section 2.1 was $\bm{\Sigma}_{ss}^{p*}=\bm{\Sigma}_{ss}^{g*}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon*}$, with stars added to denote the redundancy filtering. The reduced covariance that we use to fit the factor model is thus $\bm{\Sigma}_{ss}^{p*}-r^{-1}\bm{\Sigma}_{ss}^{\epsilon*}=\bm{\Sigma}_{ss}^{g*}=\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}$. Given the fact that maximum-likelihood estimation of $\bm{\Theta}$ is scale-free [25], we opt to estimate the loadings and uniquenesses using correlation rather than covariance matrices. Furthermore, we recognize that the sample genetic correlation matrix $\hat{\mathbf{R}}_{ss}^{g*}$ is a sufficient statistic for the population correlation matrix $\mathbf{R}_{ss}^{g*}$, meaning the maximum likelihood estimate $\hat{\bm{\Theta}}$, given a latent dimension $\tilde{m}$, can be found by minimizing the following discrepancy function [26]:
(9) $F\left[\mathbf{R}_{ss}^{g*}(\bm{\Theta});\hat{\mathbf{R}}_{ss}^{g*};\tilde{m}\right]=\ln\Big{|}\mathbf{R}_{ss}^{g*}(\bm{\Theta})\Big{|}+\operatorname{tr}\left[\hat{\mathbf{R}}_{ss}^{g*}\mathbf{R}_{ss}^{g*}(\bm{\Theta})^{-1}\right]-\ln\Big{|}\hat{\mathbf{R}}_{ss}^{g*}\Big{|}-p^{*}\text{,}$
where we constrain $\mathbf{\Lambda}^{\top}\mathbf{\Psi}^{-1}\mathbf{\Lambda}$ to be a diagonal matrix with ordered diagonal elements for identification. We replace $\hat{\mathbf{R}}_{ss}^{g*}$ by the regularized estimate $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$ as minimizing (9) requires a positive-definite and ideally well-conditioned sample genetic correlation matrix [21]. We employ a Varimax rotation once we have obtained the estimates for $\mathbf{\Lambda}$ and $\mathbf{\Psi}$ to facilitate interpretability [27]. In essence, the varimax rotation attempts to bring the loadings matrix close to having simple structure, i.e., a structure in which variables load highly on a single factor and have near zero loadings on the other factors. To avoid notational clutter we simply use $\hat{\mathbf{\Lambda}}$ to refer to either the rotated or non-rotated estimated loadings. The described factor analytic model is implemented using an efficient matrix-free estimation method [28].
2.7. Estimation of factor scores (step VI)
We would now like to obtain estimates for the factor scores $\bm{\xi}_{j}$. These factor scores are essentially estimates of the locations of genotypes in latent space, based on the many noisy but observable secondary feature measurements. They can thus be viewed as the unobserved phenotypes of a small number of latent traits that generate the observed phenotypes for the many secondary features. We will use the joint distribution of the secondary feature BLUEs and latent factors (5) in a slightly modified Thomson regression approach [29] to obtain $\hat{\bm{\xi}}_{j}$. Recalling that we estimated $\bm{\Theta}$ on the correlation scale, we first convert $\hat{\mathbf{\Lambda}}$ and $\hat{\mathbf{\Psi}}$ back to the covariance scale:
$\hat{\mathbf{\Lambda}}_{\Sigma}=\mathbf{D}\hat{\mathbf{\Lambda}}\text{ and }\hat{\mathbf{\Psi}}_{\Sigma}=\mathbf{D}\hat{\mathbf{\Psi}}\mathbf{D}\text{, where }\mathbf{D}=(\hat{\bm{\Sigma}}_{ss}^{g*}\circ\mathbf{I}_{p^{*}})^{1/2}\text{.}$
We also scale back the regularized residual correlation matrix obtained using (6) and (7):
(10) $\hat{\bm{\Sigma}}_{ss}^{\epsilon}(\vartheta_{\epsilon}^{\ddagger})=(\hat{\bm{\Sigma}}_{ss}^{\epsilon*}\circ\mathbf{I}_{p^{*}})^{1/2}\hat{\mathbf{R}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})(\hat{\bm{\Sigma}}_{ss}^{\epsilon*}\circ\mathbf{I}_{p^{*}})^{1/2}\text{.}$
We can now obtain the matrix $\mathbf{P}$ mapping the redundancy filtered secondary features to their projections on the lower-dimensional latent space through $\mathbb{E}(\bm{\xi}_{j}|\bar{\mathbf{y}}_{s(j)}^{*})$. Given that the joint distribution (5) is normal, we can obtain this conditional expectation as $\bm{\Sigma}(\bm{\Theta})_{\xi y}\left[\bm{\Sigma}(\bm{\Theta})_{yy}\right]^{-1}\bar{\mathbf{y}}_{s(j)}^{*}$ [30]. Using the Woodbury matrix identity [31], we rewrite this last expression to produce:
$\begin{split}\tilde{\bm{\xi}}_{j}&=\Big{[}\mathbf{I}_{\tilde{m}}+\hat{\mathbf{\Lambda}}^{\top}_{\Sigma}(\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*})^{-1}\hat{\mathbf{\Lambda}}_{\Sigma}\Big{]}^{-1}\hat{\mathbf{\Lambda}}^{\top}_{\Sigma}(\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*})^{-1}\bar{\mathbf{y}}_{s(j)}^{*}\\ &=\mathbf{P}^{\top}\bar{\mathbf{y}}_{s(j)}^{*}\text{.}\end{split}$
Replacing $\hat{\bm{\Sigma}}_{ss}^{\epsilon*}$ by the regularized estimate obtained using (6), (7), and (10) gives us the final expression for the estimation of the factor scores for all genotypes:
(11) $\begin{split}\tilde{\bm{\Xi}}&=\bar{\mathbf{Y}}_{s}^{*}\left[\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})\right]^{-1}\hat{\mathbf{\Lambda}}_{\Sigma}\left\{\mathbf{I}_{\tilde{m}}+\hat{\mathbf{\Lambda}}_{\Sigma}^{\top}\left[\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})\right]^{-1}\hat{\mathbf{\Lambda}}_{\Sigma}\right\}^{-1}\\ &=\bar{\mathbf{Y}}_{s}^{*}\mathbf{P}\text{.}\end{split}$
Note that $\text{var}(\bar{\mathbf{y}}_{s(j)}^{*})\rightarrow\bm{\Sigma}_{ss}^{g*}$ as $r\rightarrow\infty$, and as a result the projection $\mathbf{P}$ approaches the matrix that would be obtained using the ordinary Thomson regression approach.
In (11) we apply the projection to the BLUEs (that is, the genotypic means). In practice we project the plot-level data, however, to obtain plot level factor scores $\tilde{\bm{\Xi}}\in\mathbb{R}^{n\times\tilde{m}}$. We later use these plot-level factor scores and sums of squares to obtain estimates for the genetic and residual covariances of the latent factors and focal trait.
One of the characteristics of Thomson regression that we maintain in our modified approach is the orthogonality of the factor scores. We have thus reduced the possibly high-dimensional and collinear secondary data to a much smaller set of orthogonal factors, without losing interpretability. These orthogonal factors can easily and effectively be used in any multivariate genomic prediction approach. We opt to include (a subset of) the $\tilde{m}$ factors, the relevant covariance matrices, and the marker-based kinship matrix in the standard multivariate BLUP equations as described in the following section, thus finally integrating genomic and secondary data.
2.8. Factor selection and multivariate genomic prediction (step VII)
Having reduced the dimensionality of the secondary data, we replace the original features in the plot-level data matrix (1) by their projections:
$\mathbf{Y}_{\tilde{\Xi}f}=\begin{bmatrix}\tilde{\bm{\Xi}}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
and use sums of squares to obtain estimates for the genetic and residual covariance matrices of the factors and focal trait. We then convert the plot-level factor scores to BLUEs $\bar{\mathbf{Y}}_{\tilde{\Xi}f}$ by taking the genotypic means. Note that we have not used the focal trait data $\mathbf{y}_{f}$ in any of the previous six steps detailed above. As a result, not all of the $\tilde{m}$ factors may be relevant to the multivariate genomic prediction problem. As a form of subset selection we perform an exhaustive search for the best subset of $\tilde{m}^{\star}\leq\tilde{m}$ factors for predicting the focal trait BLUEs using the factor BLUEs in linear regression, using adjusted $R^{2}$ as the accuracy measure [32]. Following this search, we subset the BLUEs to obtain $\bar{\mathbf{Y}}_{\tilde{\Xi}^{\star}f}\in\mathbb{R}^{n\times(\tilde{m}^{\star}+1)}$ subset the covariance matrices of factors and focal trait accordingly:
$\begin{split}\hat{\bm{\Sigma}}^{g\star}&=\begin{pmatrix}\hat{\bm{\Sigma}}_{\tilde{\Xi}^{\star}\tilde{\Xi}^{\star}}^{g}&\hat{\bm{\sigma}}_{\tilde{\Xi}^{\star}f}^{g}\\ \hat{\bm{\sigma}}_{f\tilde{\Xi}^{\star}}^{g}&\hat{\sigma}_{ff}^{g}\\ \end{pmatrix}\\ \hat{\bm{\Sigma}}^{\epsilon\star}&=\begin{pmatrix}\hat{\bm{\Sigma}}_{\tilde{\Xi}^{\star}\tilde{\Xi}^{\star}}^{\epsilon}&\hat{\bm{\sigma}}_{\tilde{\Xi}^{\star}f}^{\epsilon}\\ \hat{\bm{\sigma}}_{f\tilde{\Xi}^{\star}}^{\epsilon}&\hat{\sigma}_{ff}^{\epsilon}\\ \end{pmatrix}\text{.}\end{split}$
We finally divide $\hat{\bm{\Sigma}}^{\epsilon\star}$ by the number of replicates and combine secondary and genomic data using the standard BLUP equations (see SM Section S.2).
The integration of secondary and genomic data thus follows from the use of the marker-based kinship matrix and the inclusion of the factor scores in the data matrices, along with the addition of the factors to the genetic and residual covariance matrices. If secondary data is only available for the training set (CV1), the extra features refine estimates of genetic (i.e., marker) effects for training genotypes. These improved training set estimates are then translated to test set estimates. In CV2, where secondary data is available for the test set too, the extra features provide direct information on the genetic effects for the test set, improving estimates of the focal trait directly. To evaluate the BLUP equations described in SM Section S.2 we use a fast algorithm that decreases the complexity of these calculations from $O(n_{g}^{3}(\tilde{m}+1)^{3})$ to $O(n_{g}(\tilde{m}+1)^{2}+(\tilde{m}+1)^{3})$ [33].
On a final note, consider that as the dimensionality reduction described in steps I to VI is entirely unsupervised, gfBLUP is flexible with regards to what secondary feature data is used. If secondary features are only available for the training set (CV1), estimation of the covariance matrices, redundancy filtering, regularization, estimation of $\bm{\Theta}$, etc. is based only on training data and factor scores are only obtained for the training set. In CV2, where secondary features are also available for the test set, this data can be included effortlessly as well using the multivariate BLUP equations found in SM Section S.2.
3. Simulations
We compared gfBLUP to a number of alternative approaches (siBLUP, lsBLUP, MegaLMM, deep learning (multiMLP: multivariate multilayer perceptron), univariate gBLUP, see SM Section S.5 for details on these methods) in terms of performance in different high-dimensional data settings. The following two sections provide details on the simulation process as well as the results.
3.1. Setup
[…]
Figure 2. Prediction accuracies for several methods in the CV1 and CV2 scenarios as we vary the focal trait communality, focal trait heritability, and secondary feature heritabilities. Details regarding these simulations can be found in Section 3.1.
3.2. Results
The impact of genetic parameters like the secondary feature heritabilities and focal trait communality is clear (Figure 2). With the exception of the low focal trait communality settings, the CV1 accuracy of most multivariate approaches is, at worst, on par with that of the univariate model, across the range of focal trait heritabilities. Furthermore, CV1 accuracies of all methods, except the neural networks, converge to more or less the same point ($\rho\approx 0.65$) as $h^{2}(y)$ increases, highlighting that secondary features bring the biggest benefit when the focal trait is hard to measure accurately.
Most methods significantly outperform the univariate model in CV2, provided the quality of the secondary features is high enough. Looking at the setting with the highest quality secondary features, i.e., $h^{2}(s)=0.9$ and $\bm{\lambda}_{y}\bm{\lambda}_{y}^{\top}=0.8$, improvements with respect to the univariate approach reach approximately $40\%$ to $90\%$ depending on the focal trait heritability.
Regarding the secondary feature quality, the extent to which $h^{2}(s)$ and the focal trait communality impact performance differs between methods and also seems to depend on the focal trait heritability. Consider the cases where $h^{2}(y)=0.1$, for example. In this setting the heritability of secondary features has a negligible impact on the performance of gfBLUP and MegaLMM. In the case of lsBLUP, $h^{2}(s)$ has a small impact in CV1 when the focal trait communality is high, while it is relevant in combination with all communalities in CV2. Furthermore, significant overfitting seems to occur for lsBLUP in CV2 when both $\bm{\lambda}_{y}\bm{\lambda}_{y}^{\top}$ and $h^{2}(s)$ are low. For low focal trait heritabilities, siBLUP accuracy paradoxically decreases with increasing secondary feature heritabilities, in both CV1 and CV2. This confusing behaviour of siBLUP disappears when moving to a more realistic $h^{2}(y)=0.5$. The neural network approach performs significantly worse in both CV1 and CV2, next to having extremely inconsistent performance. Overall, both MegaLMM and gfBLUP seem to capture the factor structure underlying the data to the highest degree. Both approaches perform almost identically when averaging all results. The other approaches clearly react differently to complex interactions between genetic parameters like heritabilities and communalities.
4. Analysis of CIMMYT wheat data
In addition to the extensive simulations provided in the previous section we also evaluated the different approaches in terms of accuracy and interpretability using CIMMYT wheat data [5]. The following two sections provide details on the dataset, pre-processing, and results.
4.1. Data description and pre-processing
The CIMMYT wheat dataset contains measurements on grain yield and crop canopy hyperspectral reflectivity at $62$ wavelengths between $398$nm and $847$nm. The data was collected for four breeding cycles and five managed treatments within each cycle. We limited ourselves to a subset of the data to compare gfBLUP and the other methods. The subset we analyzed (breeding cycle 2014–2015, treatment B5IR) contains $39$ trials, $1094$ unique genotypes, and reflectance measurements at ten timepoints for a total of $620$ timepoint-wavelength combinations (referred to as the secondary features). The ten timepoints span the vegetative, heading (flowering), and grain-filling growth phases. Each trial had $28$ trial-specific genotypes and two check genotypes that occurred in every trial, all replicated three times. Trials were structured according to a resolvable incomplete block design. Marker data was available for $1033$ of the $1094$ genotypes. One genotype had missing yield, leaving us with $1032$ genotypes for the analyses. We first resolved the experimental design for each secondary feature using the following linear mixed model:
(12) $y_{(i|j)kl}=\mu+g_{j}+t_{k}+r_{i(k)}+b_{l(ik)}+\epsilon_{(i|j)kl}\text{,}$
where $y_{(i|j)kl}$ is yield or any of the $620$ secondary features for replicate $i$ given genotype $j$ in trial $k$, and block $l$. The overall mean is given by $\mu$ and $g_{j}$ is the fixed genotypic effect. The random trial effect $t_{k}$ is assumed to be independently, identically, and normally distributed with mean zero and variance $\sigma_{t}^{2}$: $t_{k}\sim\mathcal{N}(0,\sigma_{t}^{2})$. Similarly, $r_{i(k)}\sim\mathcal{N}(0,\sigma_{r}^{2})$, $b_{l(ik)}\sim\mathcal{N}(0,\sigma_{b}^{2})$, and $\epsilon_{(i|j)kl}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ for the random replicate and block effects, as well as the residual effect. We then created the pseudo-CRD data by simply adding the estimated genetic effects and residuals from model (12):
$y_{q|j}=\mu+\hat{g}_{j}+\hat{\epsilon}_{(i|j)kl}\text{.}$
Note that due to the elimination of the design factors we can replace the indices $i$, $k$, and $l$ by a single index $q$ that denotes replicate $q$ of genotype $j$. We created $250$ datasets by randomly sampling $26$ trials as the training set and the remaining $13$ trials as the test set. We scaled and centered yield as well as the hyperspectral features separately for the training and test set prior to analyzing the data. We calculated the prediction accuracy as the correlation between predictions and the BLUEs ($\hat{g}_{j}$) obtained from model (12). We used a redundancy filtering threshold $\tau=0.95$ for all multivariate approaches as this provides a good trade-off between excluding redundant features and maintaining enough features to identify latent factors. Details on the methods other than gfBLUP can be found in SM Section S.5.
4.2. Results
Overall, all multivariate analyses except for deep learning performed practically the same with the CV2 predictions being more accurate than the CV1 predictions (Figure 3). None of the methods outperformed the univariate gBLUP model in CV1. Deep learning performed significantly worse than the other multivariate approaches, likely due to the relatively small dataset. Section S.8 of the SM provides results for a more practical scenario where only test set data measured during the vegetative phase is included in the CV2 predictions. Accuracies in this scenario are lower, but multivariate methods still outperform the univariate model showing that integration of genomic and phenomic data is promising even in more restrictive, realistic settings. We retained between $155$ and $172$ features after filtering using a threshold of $\tau=0.95$. The pattern of wavelengths that were retained was relatively constant across the ten dates, with most retained wavelengths representing the blue part of the spectrum. Reflectivities around 525nm, 700nm, and at 847nm are typically retained as well. A graphical overview can be found in Section S.9 of the SM.
Figure 3. Prediction accuracies for several methods in the CV1 and CV2 scenarios for the wheat yield and hyperspectral reflectivities dataset. Redundancy filtering was carried out for all multivariate analyses with a threshold $\tau=0.95$.
Figure 4. Loadings for the $62$ wavelengths and three factors found using gfBLUP and the hyperspectral data from 10-03-2015. Estimated genetic correlations between the three factors and yield are also shown.
We reanalyzed the data from one of the timepoints (10-03-2015) to assess the interpretability of gfBLUP and the other approaches. We omitted the redundancy filtering step, resulting in $62$ wavelengths. Figure 4 shows the gfBLUP factor loadings obtained through this analysis. All three factors were highly heritable at $H^{2}=0.79$ for factor $1$ and $H^{2}=0.72$ for factor $2$, and $H^{2}=0.55$ for factor $3$, where $H^{2}$ denotes the wide-sense heritability as variances were estimated using sums of squares and are thus not additive only. Genetic correlations between the factors and yield were $-0.63$ for factor $1$, $0.27$ for factor $2$, and $0.15$ for factor $3$. Through these correlations and the factor loadings, it is immediately clear that higher yielding genotypes tended to have lower reflectivities in the green-red part of the spectrum and higher reflectivities in the near-infrared (NIR) part. These results are reminiscent of the normalized difference vegetation index (NDVI) that is highly correlated with yield and is high itself for crop canopies with high NIR reflectivity and low red light reflectivity [35]. We provide details on the interpretability of the other approaches in SM Section S.6. We did not include deep learning here as it was impossible to extract any biological meaning from the parameters of the neural networks. We provide some data on the computational complexity of gfBLUP and other methods in SM Section S.7.
5. Discussion
With the recent innovations in HTP techniques and the corresponding increase in data available to plant breeders, the development of methodology for integrating this data has become increasingly relevant. We have shown that the proposed methodology gfBLUP performs on par or better than alternatives and is able to integrate secondary HTP data efficiently without loss of interpretability.
We used simulations to evaluate the impact of several genetic parameters on model performance. Of these parameters, the focal trait heritability is of particular interest. One of the key motivations for the use of secondary data in multivariate genomic prediction is the belief that this data can improve predictions in cases where the focal trait phenotype has low heritability [11]. The proposed gfBLUP performs well in such cases due to its unsupervised nature. It does not suffer from overfitting at low focal trait heritabilities like some alternative approaches.
We alse evaluated interpretability by analyzing a wheat yield and hyperspectral reflectivity dataset. We found that two of the three obtained factors were responsible for the (co)variation in reflectivities of visible and NIR light. Combined with the genetic correlations between these factors and yield we found a pattern similar to the NDVI which is known to be a good indicator of yield [36]. As large-scale data collection becomes increasingly accessible, the complexity of datasets and the importance of understanding biological implications increases as well [37]. Given the results obtained here, incorporating generative factor analysis in genomic prediction seems an effective way of tackling these challenges.
There are several options to extend the proposed methodology. While the unsupervised approach of gfBLUP worked well here, there are benefits to using a supervised approach. One way of achieving this would be to use a factor regression approach [13]. Conditioning on the prediction target may lead to higher accuracy when the focal trait training data is of sufficient quality. Another option would be to consider several ways of integrating multiple data modalities as opposed to using a single (e.g., hyperspectral) modality as we did here. There are various methods that can generally be categorized into early integration, intermediate integration, or late integration [38]. These integration approaches each try to balance parsimony and accuracy. It may be fruitful to develop methodology that finds the right balance between the parsimony and accuracy of different integration approaches within the context of genomic prediction. Finally, there is the issue of multi-year, multi-environment trials [39]. By analyzing a trait of interest measured in several environments using a multivariate model, gfBLUP could reveal latent drivers of environment-specific variety performance. Combining this approach with dynamic factor analysis [40] would provide a promising framework to model the main source of data in large breeding programs.
In any case, the gfBLUP methodology we propose here is a robust and relatively efficient approach to obtaining multivariate genomic predictions with several promising options regarding extensions and further research.
6. Data availability
Scripts to generate the simulated and hyperspectral datasets, as well as the results presented in this paper, are available at https://github.com/KillianMelsen/gfBLUP_2024. The gfBLUP methodology is implemented in an R-package available at https://github.com/KillianMelsen/gfBLUP. The hyperspectral dataset is available upon reasonable request from J. Crossa.
Supplementary Material: Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach
S.1. Notation
We denote matrices and vectors by boldface upper- and lower-case symbols, and scalars by lower-case symbols. Furthermore, we use $\bm{\Sigma}$ and $\mathbf{R}$ to denote covariance and correlation matrices. We additionally use the superscripts $p$, $g$, and $\epsilon$ to denote phenotypic, genetic, and residual covariance and correlation matrices. For example, the symbol $\bm{\Sigma}^{\epsilon}$ denotes the residual covariance matrix.
Let $\mathbf{A}\in\mathbb{R}^{n\times(p+1)}$ be some data matrix containing $p$ secondary features and a single focal trait for $n$ individuals. While we thus use $p$ for the number of secondary features as well as phenotypic covariance matrices, its meaning should be clear from the context. We assume that genotypes are replicated $r$ times following a balanced design so that the total number of individuals $n=n_{g}r$, where $n_{g}$ and $r$ represent the number of genotypes and number of replicates per genotype. Note that we assume a balanced design for notational simplicity, but that it is not a requirement. We use two indices to specify the rows of data matrices so that $\mathbf{a}_{(i|j)}$ represents the $1\times(p+1)$ data vector for replicate $i$ given genotype $j$. We use the subscripts $s$ and $f$ to denote a matrix containing only secondary features or a vector containing only the focal trait. As an example, $\mathbf{A}_{s}$ is a data matrix containing secondary features, and $\mathbf{a}_{s(i|j)}$ is the row vector of secondary features for genotype $j$, replicate $i$. The kinship matrix is denoted by $\mathbf{K}$.
We use $\circ$ and $\otimes$ to denote the Hadamard and Kronecker products. We use $diag(s_{11},\dots,s_{pp})$ to refer to a $p$-dimensional diagonal matrix with elements $s_{11},\dots,s_{pp}$. We denote the column-wise vectorization of a matrix by $\operatorname{vec}(\mathbf{A})$ and use $\operatorname{vec}^{-1}(\mathbf{A})$ for the reverse operation. The cardinality of a set $\mathcal{A}$ is denoted by $\text{card}(\mathcal{A})$. Some further information and examples on the partitioning of data and covariance matrices, as well as some general background information on genomic prediction, can be found in SM Section S.2.
S.2. General genomic prediction setting
We more or less follow the notation of Runcie and Cheng [1] and Arouisse et al. [2]. Let $\mathbf{Y}\in\mathbb{R}^{n\times(p+1)}$ be the data matrix containing unreplicated, mean-centered, and scaled measurements on $p$ secondary features and the focal trait for $n$ genotypes. We can divide the $n$ genotypes into a test and training set so that $n=n_{t}+n_{o}$ ($t$ for test and $o$ for the observed training set). The matrix $\mathbf{Y}$ can then be partitioned as:
$\mathbf{Y}=\begin{bmatrix}\mathbf{Y}_{st}&\mathbf{y}_{ft}\\ \mathbf{Y}_{so}&\mathbf{y}_{fo}\end{bmatrix}=\begin{bmatrix}\mathbf{Y}_{.t}\\ \mathbf{Y}_{.o}\end{bmatrix}\text{,}$
where $\mathbf{Y}_{st}$ and $\mathbf{Y}_{so}$ ($s$ for secondary) are the sub-matrices containing the HTP features for the test and training set, respectively. Similarly, $\mathbf{y}_{ft}$ and $\mathbf{y}_{fo}$ ($f$ for focal) are the column vectors containing the focal trait for the test and training set.
The measurements in the data matrix $\mathbf{Y}$ are composed of genetic effects, residual effects, and possibly covariates so that $\mathbf{Y}=\mathbf{X}\bm{\beta}+\mathbf{G}+\mathbf{E}$. The random matrices for the genetic and residual effects come from matrix normal distributions:
$\begin{split}\mathbf{G}&\sim\mathcal{MN}_{n,p+1}(\mathbf{0},\mathbf{K},\bm{\Sigma}^{g})\\ \mathbf{E}&\sim\mathcal{MN}_{n,p+1}(\mathbf{0},\mathbf{I}_{n},\bm{\Sigma}^{\epsilon})\text{,}\end{split}$
where the row-covariances are the kinship matrix $\mathbf{K}$ and the identity matrix $\mathbf{I}_{n}$ for the genetic and residual effects. The column-covariances are the genetic and residual covariance matrices of the secondary features and focal trait. Assuming that $\mathbf{Y}$ is only composed of genetic and residual effects, a standard result shows that the column-wise vectorization of $\mathbf{Y}$, denoted as $\operatorname{vec}(\mathbf{Y})$, follows a multivariate normal distribution:
$\operatorname{vec}(\mathbf{Y})=\operatorname{vec}(\mathbf{G})+\operatorname{vec}(\mathbf{E})\sim\mathcal{N}_{n(p+1)}(\mathbf{0},\bm{\Sigma}^{g}\otimes\mathbf{K}+\bm{\Sigma}^{\epsilon}\otimes\mathbf{I}_{n})\text{,}$
where $\otimes$ denotes the Kronecker product. Similarly to the data matrix itself, we can partition the kinship and covariance matrices as well:
$\mathbf{K}=\begin{bmatrix}\mathbf{K}_{tt}&\mathbf{K}_{to}\\ \mathbf{K}_{ot}&\mathbf{K}_{oo}\end{bmatrix}\text{,}$
where $\mathbf{K}_{tt}$ and $\mathbf{K}_{oo}$ contain the genetic relatedness coefficients within the test and training sets, respectively. The genetic relatedness coefficients between test and training genotypes are given by $\mathbf{K}_{to}=(\mathbf{K}_{ot})^{\top}$. Furthermore,
$\begin{split}\bm{\Sigma}^{g}&=\begin{bmatrix}\bm{\Sigma}_{ss}^{g}&\bm{\sigma}_{sf}^{g}\\ \bm{\sigma}_{fs}^{g}&\sigma_{ff}^{g}\\ \end{bmatrix}=\begin{bmatrix}\bm{\sigma}_{s.}^{g}\\ \bm{\sigma}_{f.}^{g}\end{bmatrix}\\ \bm{\Sigma}^{\epsilon}&=\begin{bmatrix}\bm{\Sigma}_{ss}^{\epsilon}&\bm{\sigma}_{sf}^{\epsilon}\\ \bm{\sigma}_{fs}^{\epsilon}&\sigma_{ff}^{\epsilon}\\ \end{bmatrix}=\begin{bmatrix}\bm{\sigma}_{s.}^{\epsilon}\\ \bm{\sigma}_{f.}^{\epsilon}\end{bmatrix}\text{,}\end{split}$
where $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$ contain the genetic and residual covariances among secondary features, $\sigma_{ff}^{g}$ and $\sigma_{ff}^{\epsilon}$ are the scalar focal trait genetic and residual variances, and $\bm{\sigma}_{sf}^{g}=(\bm{\sigma}_{fs}^{g})^{\top}$ and $\bm{\sigma}_{sf}^{\epsilon}=(\bm{\sigma}_{fs}^{\epsilon})^{\top}$ contain the genetic and residual covariances between secondary features and the focal trait.
As mentioned above, there are generally speaking three genomic prediction scenarios depending on whether secondary data is available for just the training set (CV1), both training and test set (CV2), or not at all (univariate genomic prediction). We now review the expressions for the training ($\mathbf{G}_{o}$) and test set ($\mathbf{G}_{t}$) BLUPs in these three scenarios.
In the univariate scenario, we only use information on the focal trait of the training set ($\mathbf{y}_{fo}$) to obtain predictions for the test set:
(S1) $\begin{split}\hat{\mathbf{V}}&=\hat{\sigma}_{ff}^{g}\otimes\mathbf{K}_{oo}+\hat{\sigma}_{ff}^{\epsilon}\otimes\mathbf{I}_{n_{o}}\\ \hat{\mathbf{g}}^{(uni)}_{fo}&=\left(\hat{\sigma}_{ff}^{g}\otimes\mathbf{K}_{oo}\right)\hat{\mathbf{V}}^{-1}\left(\mathbf{y}_{fo}-\mathbf{X}_{o}\hat{\bm{\beta}_{f}}\right)\\ \hat{\mathbf{g}}^{(uni)}_{ft}&=\mathbb{E}\Bigl{(}\mathbf{g}_{ft}|\mathbf{y}_{fo}\Bigr{)}\\ &=\left(\hat{\sigma}_{ff}^{g}\otimes\mathbf{K}_{to}\right)\hat{\mathbf{V}}^{-1}\left(\mathbf{y}_{fo}-\mathbf{X}_{o}\hat{\bm{\beta}_{f}}\right)\\ &=\mathbf{K}_{to}\mathbf{K}_{oo}^{-1}\hat{\mathbf{g}}^{(uni)}_{fo}\text{,}\end{split}$
where $\hat{\mathbf{g}}^{(uni)}_{fo}$ and $\hat{\mathbf{g}}^{(uni)}_{ft}$ are the focal trait univariate BLUPs for the training and test set. Note that covariates can be included in the term $\mathbf{X}_{o}\hat{\bm{\beta}_{f}}$. Also note that as $\hat{\sigma}_{ff}^{g}$ and $\hat{\sigma}_{ff}^{\epsilon}$ are scalars, $\hat{\mathbf{V}}=\hat{\sigma}_{ff}^{g}\mathbf{K}_{oo}+\hat{\sigma}_{ff}^{\epsilon}\mathbf{I}_{n_{o}}$.
In the multivariate CV1 scenario where secondary features are only available for the training set, predictions can be obtained using:
(S2) $\begin{split}\hat{\mathbf{V}}&=\hat{\bm{\Sigma}}^{g}\otimes\mathbf{K}_{oo}+\hat{\bm{\Sigma}}^{\epsilon}\otimes\mathbf{I}_{n_{o}}\\ \hat{\mathbf{g}}^{(CV1)}_{fo}&=\left(\hat{\bm{\sigma}}^{g}_{f.}\otimes\mathbf{K}_{oo}\right)\hat{\mathbf{V}}^{-1}\operatorname{vec}\left(\mathbf{Y}_{.o}-\mathbf{X}_{o}\hat{\bm{\beta}}\right)\\ \hat{\mathbf{g}}^{(CV1)}_{ft}&=\mathbb{E}\Bigl{(}\mathbf{g}_{ft}|\mathbf{Y}_{.o}\Bigr{)}\\ &=\left(\hat{\bm{\sigma}}^{g}_{f.}\otimes\mathbf{K}_{to}\right)\hat{\mathbf{V}}^{-1}\operatorname{vec}\left(\hat{\mathbf{Y}}_{.o}-\mathbf{X}_{o}\hat{\bm{\beta}}\right)\\ &=\mathbf{K}_{to}\mathbf{K}_{oo}^{-1}\hat{\mathbf{g}}^{(CV1)}_{fo}\text{,}\end{split}$
where $\mathbf{Y}_{.o}-\mathbf{X}_{o}\hat{\bm{\beta}}$ is vectorized. Expression (S2) can be evaluated using a fast algorithm that decreases the time complexity for the calculation of the focal trait training set BLUPs $\hat{\mathbf{g}}^{(CV1)}_{fo}$ from $O(n_{o}^{3}p^{3})$ to $O(n_{o}p^{2}+p^{3})$ by avoiding the naive inversion of $\hat{\mathbf{V}}$ [3].
The CV2 multivariate predictions can be obtained directly [2] or through a two-step approach [1]. We prefer the two-step approach because the fast algorithm can be used to obtain all CV1 training set BLUPs, $\hat{\mathbf{G}}^{(CV1)}_{.o}$:
(S3) $\begin{split}\hat{\mathbf{V}}&=\hat{\bm{\Sigma}}^{g}\otimes\mathbf{K}_{oo}+\hat{\bm{\Sigma}}^{\epsilon}\otimes\mathbf{I}_{n_{o}}\\ \hat{\mathbf{G}}^{(CV1)}_{.o}&=\operatorname{vec}^{-1}\left[\left(\hat{\bm{\Sigma}}^{g}\otimes\mathbf{K}_{oo}\right)\hat{\mathbf{V}}^{-1}\operatorname{vec}\left(\mathbf{Y}_{.o}-\mathbf{X}_{o}\hat{\bm{\beta}}\right)\right]\text{.}\end{split}$
We then convert these training set BLUPs to intermediate CV1 test set BLUPs and combine those with the observed secondary features for the test set to finally obtain CV2 BLUPs for the test set focal trait:
(S4) $\begin{split}\hat{\mathbf{V}}&=\hat{\bm{\Sigma}}_{ss}^{g}\otimes\mathbf{K}_{tt}^{-1}+\hat{\bm{\Sigma}}_{ss}^{\epsilon}\otimes\mathbf{I}_{n_{t}}\\ \hat{\mathbf{g}}^{(CV2)}_{ft}&=\mathbb{E}\Bigl{(}\mathbf{g}_{ft}|\mathbf{Y}_{st},\mathbf{Y}_{.o}\Bigr{)}\\ &=\mathbf{K}_{to}\mathbf{K}_{oo}^{-1}\hat{\mathbf{g}}^{(CV1)}_{fo}+\left(\hat{\bm{\sigma}}^{g}_{fs}\otimes\mathbf{K}_{tt}^{-1}\right)\hat{\mathbf{V}}^{-1}\operatorname{vec}\left(\mathbf{Y}_{st}-\mathbf{K}_{to}\mathbf{K}_{oo}^{-1}\hat{\mathbf{G}}^{(CV1)}_{so}\right)\text{.}\end{split}$
In all cases, $\bm{\Sigma}^{g}$ and $\bm{\Sigma}^{\epsilon}$ must be estimated from the data.
S.3. Variance of BLUEs
Consider the secondary feature BLUEs for genotype $j$ which are simply the genotypic means in the case of a completely randomized experimental design:
$\bar{\mathbf{y}}_{s(j)}=\mathbf{g}_{s(j)}+\bar{\bm{\epsilon}}_{s(j)}=\mathbf{g}_{s(j)}+\dfrac{1}{r}\sum_{i=1}^{r}\bm{\epsilon}_{s(i|j)}\text{,}$
where $\mathbf{g}_{s(j)}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{ss}^{g})$ and $\bm{\epsilon}_{s(i|j)}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{ss}^{\epsilon})$. All $r$ residual vectors for the replicates of genotype $j$ are assumed iid, meaning $\text{var}(\bm{\epsilon}_{s(i|j)})=\text{var}(\bm{\epsilon}_{s(i^{\prime}|j)})=\bm{\Sigma}_{ss}^{\epsilon}$ and the cross-covariance matrix for any two replicates of genotype $j$, $\text{cov}(\bm{\epsilon}_{s(i|j)},\bm{\epsilon}_{s(i^{\prime}|j)})$, is a zero matrix. As a result,
$\text{var}\left(\sum_{i=1}^{r}\bm{\epsilon}_{s(i|j)}\right)=\sum_{i=1}^{r}\text{var}\left(\bm{\epsilon}_{s(i|j)}\right)=r\bm{\Sigma}_{ss}^{\epsilon}\text{.}$
Now we divide the sum of the $r$ residual vectors for genotype $j$ by the number of replicates $r$ to obtain the vector of mean residuals for genotype $j$:
$\bar{\bm{\epsilon}}_{s(j)}=\dfrac{1}{r}\sum_{i=1}^{r}\bm{\epsilon}_{s(i|j)}\text{.}$
Using the fact that $\text{var}(a\mathbf{x})=a^{2}\text{var}(\mathbf{x})$ for some scalar $a$ and random vector $\mathbf{x}$, we find that
$\text{var}\left(\bar{\bm{\epsilon}}_{s(j)}\right)=\left(\dfrac{1}{r}\right)^{2}\text{var}\left(\sum_{i=1}^{r}\bm{\epsilon}_{s(i|j)}\right)=r^{-2}r\bm{\Sigma}_{ss}^{\epsilon}=r^{-1}\bm{\Sigma}_{ss}^{\epsilon}\text{.}$
Combined with the assumptions that $\mathbf{g}_{s(j)}\ConIndepNat\bm{\epsilon}_{s(i|j)}$, and thus that $\mathbf{g}_{s(j)}\ConIndepNat\bar{\bm{\epsilon}}_{s(j)}$, we obtain the covariance of the BLUEs as
$\text{var}\left(\bar{\mathbf{y}}_{s(j)}\right)=\text{var}\left(\mathbf{g}_{s(j)}\right)+\text{var}\left(\bar{\bm{\epsilon}}_{s(j)}\right)=\bm{\Sigma}_{ss}^{g}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon}\text{.}$
S.4. Efficient penalty optimization
We use Brent’s algorithm [11] implemented in the optimize() function in R to find the optimal values for $\vartheta_{g}$ and $\vartheta_{\epsilon}$ as described in the main text. In a naive implementation we obtain the mean cross-validated negative log-likelihood at each iteration as:
(S5) $\varphi^{K}(\vartheta)=\dfrac{1}{K}\sum_{k=1}^{K}w_{k}\left\{\text{ln}\Big{|}\Bigl{(}\mathbf{R}(\vartheta)\Bigr{)}_{\neg k}\Big{|}+\operatorname{tr}\left[\mathbf{R}_{k}\Bigl{(}\mathbf{R}(\vartheta)\Bigr{)}_{\neg k}^{-1}\right]\right\}\text{,}$
where we have omitted some sub- and superscripts for clarity. Note that $(\mathbf{R}(\vartheta))_{\neg k}\in\mathbb{R}^{p\times p}$ is the penalized genetic or residual correlation matrix given the corresponding penalty $\vartheta$ at the current iteration. Naive minimization thus requires inverting $(\mathbf{R}(\vartheta))_{\neg k}=\left(1-\vartheta\right)\mathbf{R}_{\neg k}+\vartheta\mathbf{I}$ at each iteration. Note that
$\Bigl{[}\mathbf{R}(\vartheta)\Bigr{]}_{\neg k}^{-1}=\mathbf{U}\Bigl{[}\mathbf{\Lambda}(\vartheta)\Bigr{]}^{-1}\mathbf{U}^{\top}\text{,}$
where $\mathbf{U}$ contains the eigenvectors of the unregularized correlation matrix $\mathbf{R}_{\neg k}$ and
$\mathbf{\Lambda}(\vartheta)=\left(1-\vartheta\right)\mathbf{\Lambda}+\vartheta\mathbf{I}_{p}$
with $\mathbf{\Lambda}$ a diagonal matrix containing the eigenvalues of $\mathbf{R}_{\neg k}$. We now rewrite (S5) as:
(S6) $\varphi^{K}(\vartheta)=\dfrac{1}{K}\sum_{k=1}^{K}w_{k}\left\{\text{ln}\prod_{i=1}^{p}\mathbf{\Lambda}(\vartheta)_{ii}+\operatorname{tr}\left[\mathbf{R}_{k}\mathbf{U}\Bigl{(}\mathbf{\Lambda}(\vartheta)\Bigr{)}^{-1}\mathbf{U}^{\top}\right]\right\}\text{.}$
We can then use the invariance of the trace operator to cyclic permutations to rewrite (S6) as:
(S7) $\varphi^{K}(\vartheta)=\dfrac{1}{K}\sum_{k=1}^{K}w_{k}\left\{\text{ln}\prod_{i=1}^{p}\mathbf{\Lambda}(\vartheta)_{ii}+\operatorname{tr}\left[\Bigl{(}\mathbf{\Lambda}(\vartheta)\Bigr{)}^{-1}\mathbf{A}\right]\right\}\text{,}$
where $\mathbf{A}=\mathbf{U}^{\top}\mathbf{R}_{k}\mathbf{U}$. Within the trace operation in (S7) we are now calculating the matrix product of the diagonal matrix $\left(\mathbf{\Lambda}(\vartheta)\right)^{-1}$ and the non-diagonal matrix $\mathbf{A}$, of which the latter is independent of $\vartheta$. This can be rewritten as a summation of divisions:
(S8) $\varphi^{K}(\vartheta)=\dfrac{1}{K}\sum_{k=1}^{K}w_{k}\left[\text{ln}\prod_{i=1}^{p}\mathbf{\Lambda}(\vartheta)_{ii}+\sum_{j=1}^{p}\dfrac{\mathbf{A}_{jj}}{\mathbf{\Lambda}(\vartheta)_{jj}}\right]\text{.}$
We can calculate the eigenvalues $\mathbf{\Lambda}$ as well as the diagonal entries of $\mathbf{A}=\mathbf{U}^{\top}\mathbf{R}_{k}\mathbf{U}$ before starting the optimization. As a result, each iteration scales linearly with $p$ resulting in faster optimization, especially if the optimal penalty values are near the bounds.
S.5. Alternative methods
S.5.1. Univariate gBLUP
The baseline model that we compare all multivariate approaches to is the standard univariate gBLUP model:
$\mathbf{y}=\mathbf{Z}\mathbf{u}+\bm{\epsilon}\text{,}$
where $\mathbf{y}$ is the $n\times 1$ vector containing the mean-centered and scaled focal trait phenotypes, $\mathbf{Z}$ is the incidence matrix linking individuals to the BLUPs in $\mathbf{u}\sim\mathcal{N}(\mathbf{0},\sigma_{g}^{2}\mathbf{K})$, and $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$ contains the residuals. We used the R-package rrBLUP version 4.6.1 [4] to fit univariate gBLUP models (see S1). The marker-based kinship matrix was created using the R-package statgenGWAS version 1.0.9.
S.5.2. siBLUP
The first alternative multivariate approach that we included is a modified version of the bivariate gBLUP model using a regularized selection index as proposed by Lopez-Cruz et al. [5] and also implemented by Arouisse et al. [2]. In short, siBLUP uses a regularized phenotypic covariance matrix of the secondary features and their genetic covariances with the focal trait to obtain a weight for every secondary feature. We use these weights to calculate a single linear combination of the secondary features for each individual termed the (regularized) selection index. We then use this index as the single secondary feature in the multivariate gBLUP equations (SM Section S.2).
Like gfBLUP, the modified version of siBLUP takes data containing replicates from a completely randomized design. Next, we obtain $\hat{\bm{\Sigma}}_{ss}^{g}$ and $\hat{\bm{\Sigma}}_{ss}^{\epsilon}$ using sums of squares, after which we redundancy filter the dataset so that no two of the remaining secondary features have an absolute genetic correlation $|\rho^{g}|$ equal to, or exceeding a threshold $\tau$.
We then convert the redundancy filtered, plot-level phenotypic covariance matrix $\hat{\bm{\Sigma}}_{ss}^{p*}=\hat{\bm{\Sigma}}_{ss}^{g*}+\hat{\bm{\Sigma}}_{ss}^{\epsilon*}$ of the remaining secondary features to a correlation matrix which is regularized:
$\hat{\mathbf{R}}_{ss}^{p*}(\vartheta_{p})=\left(1-\vartheta_{p}\right)\hat{\mathbf{R}}_{ss}^{p*}+\vartheta_{p}\mathbf{I}_{p^{*}}\text{,}$
where the optimal penalty $\vartheta^{\ddagger}_{p}$ is found using 5-fold cross validation as described in the main text using the number of individuals in each fold as the weights $w_{k}$. We then reconvert the resulting regularized phenotypic correlation matrix to the covariance matrix $\hat{\bm{\Sigma}}_{ss}^{p*}(\vartheta^{\ddagger}_{p})$ and use it to calculate the vector of selection index weights $\hat{\bm{\gamma}}^{SI}$.
To calculate the selection index weights, we first obtain the $p^{*}\times 1$ column vector $\hat{\bm{\sigma}}_{sf}^{g*}$ containing the genetic covariances between the remaining secondary features and the focal trait using sums of squares. Next, we calculate the vector of selection index weights $\hat{\bm{\gamma}}^{SI}$:
$\hat{\bm{\gamma}}^{SI}=\left[\hat{\bm{\Sigma}}_{ss}^{p*}(\vartheta^{\ddagger}_{p})\right]^{-1}\hat{\bm{\sigma}}_{sf}^{g*}\text{,}$
and use it to calculate the plot-level selection indices:
$\hat{\mathbf{y}}_{SI}=\mathbf{Y}_{s}^{*}\hat{\bm{\gamma}}^{SI}\text{,}$
where $\hat{\mathbf{y}}_{SI}$ is the $n\times 1$ column vector containing the selection index for each individual. The $n\times p^{*}$ matrix $\mathbf{Y}_{s}^{*}$ contains the original (redundancy filtered) secondary features.
We can now replace the original secondary features in the data matrix $\mathbf{Y}_{sf}$ by the selection indices:
$\mathbf{Y}_{SI,f}=\begin{bmatrix}\hat{\mathbf{y}}_{SI}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
followed by using sums of squares to obtain $\hat{\bm{\Sigma}}^{g}$ and $\hat{\bm{\Sigma}}^{\epsilon}$ which give us the genetic and residual covariances between the selection index and the focal trait. We then reduce the data to BLUEs $\bar{\mathbf{Y}}_{SI,y}$ by calculating the genoypic means, divide $\hat{\bm{\Sigma}}^{\epsilon}$ by the number of replicates, and combine the data, covariance matrices, and the kinship $\mathbf{K}$ in the BLUP equations (see SM Section S.2) to obtain test set predictions.
S.5.3. lsBLUP
lsBLUP also reduces the secondary features to a single dimension [2]. For lsBLUP, this secondary feature is a LASSO regression prediction of the focal trait using the redundancy filtered set of secondary features.
We use sums of squares obtained from replicated completely randomized design data to estimate $\bm{\Sigma}_{ss}^{g}$. We then use this genetic covariance matrix of the original secondary features (converted to the correlation scale) to redundancy filter the data. After filtering, the LASSO regression coefficients $\hat{\bm{\beta}}^{LS}$ are obtained using the R-package glmnet version 4.1.6 [6]. To do so, we first reduce the redundancy filtered data to genotypic means. We then use the cv.glmnet() function to determine the optimal (i.e., minimum mean cross-validated error) LASSO penalty $\lambda^{\ddagger}$ using 5-fold cross-validation, thereby obtaining the vector of coefficients given this optimal penalty $\hat{\bm{\beta}}^{LS}(\lambda^{\ddagger})$.
If all coefficients have been regularized to 0, lsBLUP returns simple univariate gBLUP predictions for the test set using the rrBLUP R-package [4]. If not, we calculate plot-level LASSO predictions:
$\hat{\mathbf{y}}_{LS}=\mathbf{Y}_{s}^{*}\hat{\bm{\beta}}^{LS}(\lambda^{\ddagger})\text{,}$
where $\mathbf{Y}_{s}^{*}$ is the matrix containing the redundancy filtered observed secondary features. We then replace the original secondary features in the data matrix $\mathbf{Y}_{sf}$ by the LASSO predictions:
$\mathbf{Y}_{LS,y}=\begin{bmatrix}\hat{\mathbf{y}}_{LS}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
and compute the CV1 or CV2 test set focal trait BLUPs as described for siBLUP and gfBLUP, using $\hat{\mathbf{y}}_{LSP}$ as the single secondary feature.
S.5.4. MegaLMM
MegaLMM was introduced as a method to make genomic predictions using many secondary features [7]. In essence, it decomposes the $p$ original features into two parts. One part is projected onto $M$ latent factors while the other part represents the $p$ independent errors for each secondary feature. Independent linear mixed models are then fit for these $M+p$ decorrelated traits. For the full details on MegaLMM see Runcie et al. [7].
To be able to make consistent comparisons, we redundancy filtered the full set of secondary features as we did for other methods prior to calculating the BLUEs and running MegaLMM. We initialized and ran MegaLMM with default settings except for the following parameters: for the MegaLMM_control() function drop0_tol was set to $10^{-10}$, h2_divisions was set to $20$, burn was set to 00 (we did not use the built-in function to automatically discard burn-in iterations), and thin was set to $2$. The number of modeled latent factors $M$ (argument K) was set to $16$ for the simulated data and $20$ for the hyperspectral data. For the priors we again mostly used the default settings with two exceptions. For tot_Y_var, we set nu to $10$ for both simulated and hyperspectral data. We set nu for tot_F_var to $100{,}000$ and $20$ for the simulated and hyperspectral data, respectively.
After initializing MegaLMM we first ran $1000$ burn-in iterations, reordering the factors after each $100$ iterations, for both simulated and hyperspectral data. We then ran another 2000 iterations for the simulated data and 500 iterations for the hyperspectral data. Posterior predictions for each genotype were calculated at each iteration, with the final predictions being equal to the posterior means. We used version 0.1.0 of the MegaLMM R-package [8].
S.5.5. Deep learning
Deep learning provides a flexible, non-parametric alternative to the models described above. Here, we limited ourselves to CV1 and CV2 versions of a simple multilayer perceptron (multiMLP). The only difference between these versions is the inclusion of the scaled and centered, redundancy filtered set of secondary features in the output layer (CV1) or input layer (CV2).
We considered the number of hidden layers, number of neurons per hidden layer, and the dropout regularization rate as the hyperparameters to be tuned separately for each (simulated) dataset. Other hyperparameters such as the learning rate, early stopping patience, optimizer, etc. were fixed. After an initial exploration of possible hyperparameter values, we decided to use either $4$ or $8$ hidden layers, with $16$ or $64$ neurons each. The dropout rate was set to either $0.1$ or $0.3$. In addition to these values, we opted for the root mean squared propagation optimizer (RMSProp), a learning rate of $5\times 10^{-4}$, the rectified linear unit (ReLU) activation function in all relevant layers, a batch size of $128$, and mean squared error (mse) for the loss function.
To obtain predictions for each dataset, we first redundancy filtered the secondary features like in the other multivariate approaches, followed by reducing the data to BLUEs. We then divided the training individuals over five folds to determine the within-training set mean cross validated prediction accuracy for the eight different hyperparameter combinations. We trained each of the eight networks five times (once for each fold) for at most $25$ epochs. Training could be stopped early based on the loss for a $10\%$ validation split using a patience of $10$ epochs. We then determined the combination of hyperparameter values giving the highest mean prediction accuracy to be used for training the final network on the full training set. This was done for at most $100$ epochs using early stopping with a patience of $20$ epochs and a $10\%$ validation split. We then obtained test set predictions by feeding the SNP marker matrix (and filtered secondary features for the CV2 multiMLP) for the test set into the trained networks.
We used the keras (v2.13.0) and tensorflow (v2.15.0.9) R packages as an interface to Python (v3.10) and TensorFlow (v2.13.1) to implement the deep learning [9, 10].
S.6. Interpretability
We have stressed the great ease of interpreting gfBLUP’s model parameters in terms of biological relevance. Here we present some additional results with regards to interpretability of the approaches that we compared gfBLUP to. We ran si-, lsBLUP, and MegaLMM on the non-redundancy filtered hyperspectral data from 10-03-2015 as we did for gfBLUP in Section 4.2. We then looked at the relevant model parameters that included the selection index weights for siBLUP, LASSO coefficients for lsBLUP, and (mean) posterior loadings for MegaLMM.
Figure S1. LASSO regression coefficients (A) and selection index weights (B) for the $62$ wavelengths obtained from running lsBLUP and siBLUP on the hyperspectral data from 10-03-2015. Estimated genetic correlations between the LASSO prediction (LS) or selection index (SI) and yield (Y) are also shown.
Figure S2. MegaLMM mean posterior loadings for the $62$ wavelengths on the $5$ (A) and $10$ (B) factors used to run MegaLMM on the hyperspectral data from 10-03-2015. Loadings of yield on the factors are also shown.
Starting with lsBLUP and siBLUP (Figure S1) we immediately recognize the difficulties associated with compressing the collinear data to a single dimension. In the case of lsBLUP we see sign-flipping in the coefficients around the $700$nm wavelength. Reflectivities between $650$nm and $700$nm show an increasingly positive relationship with yield following from the genetic correlation of $0.81$ between the LASSO prediction and yield. After $700$nm the relationship rapidly becomes negative, before moving towards 00 again.
The selection index weights of siBLUP are difficult to interpret as well. The relationship between reflectivity and yield follows a somewhat similar pattern to lsBLUP, given the highly positive genetic correlation between the selection index and yield ($\rho^{g}=0.81$). The relationship becomes increasingly positive, followed by a strong decrease after $700$nm, before finally increasing again. The relationship then continues to oscillate in the remainder of the infra-red part of the spectrum.
While si- and lsBLUP do not require any hyperparameter tuning, their lack of flexibility with regards to the selection of an appropriate number of dimensions clearly limits their interpretability in the case of more complex data.
MegaLMM does provide the flexibility of choosing the number of factors ($M$) and thereby the dimensionality, but leaves this choice to the user. As a result, the ease of interpretation heavily relies on an appropriate choice of $M$, which may not be easy to make without any prior knowledge (Figure S2). We opted for values of $M=5$ and $M=10$, which seemed appropriate given the fact we found $3$ factors using gfBLUP. Note that while the main factors representing the infrared and visible part of the spectrum are still present (F5 and F2 for $M=5$, F3 and F1 for $M=10$), ease of interpretation goes down as the number of factors increases. This is mainly due to the loadings of yield, which are analogous to the genetic correlations between yield and factors in gfBLUP, becoming divided over several relevant factors as well as non-relevant factors.
Furthermore, we observe strong non-stationary behavior in the loading MCMC chains for both $M=5$ (Figure S3) and $M=10$ (Figure S4), even after $10{,}000$ burn-in and a further $100{,}000$ posterior iterations for wavelengths around $700$nm (thinning rate of $2$).
Figure S3. MegaLMM ($5$ factors, $100{,}000$ iterations with thinning rate of $2$ after $10{,}000$ burn-in iterations) traceplots for the posterior samples for loadings of wavelengths around 700 nm on the $5$ factors (top to bottom).
Figure S4. MegaLMM ($10$ factors, $100{,}000$ iterations with thinning rate of $2$ after $10{,}000$ burn-in iterations) traceplots for the posterior samples for loadings of wavelengths around 700 nm on factor $1$ $5$. (Continues on next page).
Figure S4. (Continued). MegaLMM ($10$ factors, $100{,}000$ iterations with thinning rate of $2$ after $10{,}000$ burn-in iterations) traceplots for the posterior samples for loadings of wavelengths around 700 nm on factor $6$ $10$.
S.7. Computational complexity
Obtaining test set predictions for a single CV2 hyperspectral dataset (including all $620$ wavelength-date features) took approximately $1.53$ seconds for the univariate gBLUP model, $7.83$ seconds for lsBLUP, $218$ seconds for siBLUP, $43.3$ seconds for gfBLUP, $1030$ seconds for MegaLMM, and $93.2$ seconds for the CV2 multivariate MLP.
We also simulated datasets containing $p=100$ to $p=2500$ secondary features representing $10$ underlying factors. We analyzed these datasets using gfBLUP, obtaining five runtime measurements (after a single warmup run) for each step in the gfBLUP prediction pipeline at each value of $p$, showing that the regularization dominates the pipeline’s total runtime throughout the range of $p$ (Figure S5).
All computational times were obtained on Windows 11 Pro 23H2, R v4.2.2, using an Intel i5-13600KF CPU. Runtimes for deep learning were obtained using TensorFlow v2.15.1 and Python v3.10 through the keras v2.15.0 and tensorflow v2.16.0 R packages.
Figure S5. Composition of total runtime of gfBLUP at different numbers of secondary features $p$. The red line shows the total runtime in seconds.
S.8. CV2 using vegetative data
From a practical point of view, including test set secondary data measured during the full crop growth cycle (including grain filling) in CV2 single trial genomic prediction makes little sense. A more relevant scenario is to only include test set secondary features measured during the vegetative phase. Plots in field trials can then be culled based on multivariate genomic predictions to avoid crossing and harvesting of unpromising genotypes, saving a significant amount of resources.
Figure S6. Prediction accuracies for several methods in the CV1, CV2, and CV2VEG (secondary features measured during the vegetative phase only) scenarios for the wheat yield and hyperspectral reflectivities dataset. Redundancy filtering was carried out for all multivariate analyses with a threshold $\tau=0.95$.
Figure S6 provides results for such a scenario. While CV2 predictions using features from the vegetative phase only (CV2VEG) are lower than those obtained when including all data (CV2), multivariate approaches still outperform the univariate model by a significant margin.
S.9. Redundancy filtering
Figure S7 provides a graphical overview of the features that remained after filtering at a threshold of $\tau=0.95$.
Figure S7. Features (wavelength-date combinations) that remain after redundancy filtering using a threshold of $\tau=0.95$.